OpenStudy (anonymous):

(6x^2)^4 Can someone help me simplify this?

OpenStudy (anonymous):

1296x^8

OpenStudy (ash2326):

We have \[(6x^2)^4\] let's open the bracket \[6^4 \times (x^2)^4\] we know 6^4=1296 and (x^2)^4=x^8, so we get \[1296x^8\]

OpenStudy (ash2326):

Brad would you like me to walk you through Open Study?

OpenStudy (anonymous):

what else can i do on this?

OpenStudy (anonymous):

i just need to post two more questions for now :)

OpenStudy (ash2326):

Yeah, just go to the math chat below. We'll talk there. Post your questions, I'll try to help you:)

OpenStudy (anonymous):

wait where is math chat?

OpenStudy (ash2326):

On the bottom of the page , you'd see mathematics chat, which is continuously changing color to orange and green , click on it and type your message

OpenStudy (anonymous):

sorry i dont see it

OpenStudy (ash2326):

Just click on Open Study on the top left. You'd come to home page and then you'll find maths chat below